OpenStudy (anonymous):

\[\sum _{n=1}^{\infty } \frac{1}{n^2}\] What's the Sum

OpenStudy (anonymous):

F(x)=|x| from -L to L May be represented using Fourier series as \[|x|=\frac{L}{2}- \frac{(4L)}{\pi ^2}\sum _{n=1,3,5}^{\infty } \frac{1}{n^2}\text{Cos}\left[\frac{\text{n$\pi $} x}{L}\right]\] \[F[0]=\frac{L}{2}- \frac{(4L)}{\pi ^2}\sum _{n=1,3,5}^{\infty } \frac{1}{n^2}\] \[0=\frac{1}{2}- \frac{(4)}{\pi ^2}\sum _{n=1,3,5}^{\infty } \frac{1}{n^2}\] \[\frac{1}{2}= \frac{(4)}{\pi ^2}\sum _{n=1,3,5}^{\infty } \frac{1}{n^2}\] \[\frac{\pi ^2}{8}=\left(1+\frac{1}{3^2}+\frac{1}{5^2}+\frac{1}{7^2}+\text{...}.\right)\] \[\sum _{n=1}^{\infty } \frac{1}{n^2}=\left(1+\frac{1}{3^2}+\frac{1}{5^2}+\frac{1}{7^2}+\text{...}.\right)+\left(\frac{1}{2^2}+\frac{1}{4^2}+\frac{1}{6^2}+\text{...}.\right)\] \[\sum _{n=1}^{\infty } \frac{1}{n^2}=\left(1+\frac{1}{3^2}+\frac{1}{5^2}+\frac{1}{7^2}+\text{...}.\right)+\frac{1}{4}\left(\frac{1}{1^2}+\frac{1}{2^2}+\frac{1}{3^2}+\frac{1}{4^2}+\text{...}.\right)\] \[\frac{3}{4}\left(\frac{1}{1^2}+\frac{1}{2^2}+\frac{1}{3^2}+\frac{1}{4^2}+\text{...}.\right)=\frac{\pi ^2}{8}\] \[\frac{3}{4}\left(\frac{1}{1^2}+\frac{1}{2^2}+\frac{1}{3^2}+\frac{1}{4^2}+\text{...}.\right)=\frac{\pi ^2}{6}\]

OpenStudy (jamesj):

Yes. That's one proof, and a nice one too. This is famous problem and there quite a few proofs floating around.

OpenStudy (anonymous):

Is there a Proof which can be understood by me?

OpenStudy (anonymous):

I think Euler did something to proof this as well?

OpenStudy (anonymous):

Well it's just algebra after the fourier series part

OpenStudy (anonymous):

I don't have a Clue about the fourier series I always end up readin furier as fuhrer

OpenStudy (anonymous):

Oh I see, I think maybe I got it; little bit of fourier series.

OpenStudy (anonymous):

What gave you the intuition to use Fourier Series? Or is that just somewhat of a standard trick for this series?